# Bioverse: GMT and ELT Direct Imaging and High-Resolution Spectroscopy Assessment – Surveying Exo-Earth O2 and Testing the Habitable Zone Oxygen Hypothesis

### v1

@@ 2.3 Time Needed to Probe O2
With our new instrument model, we now have to determine how long it will take to probe O2. For each simulated exoplanet, we first computed the total number of hours the planet would be observable from the GMT and ELT locations for one year of observations ($N_{\text{hr-obs}}$) accounting for the aforementioned ground-based observing limitations and exoplanet orbital properties. Since the modeled planets change illumination and scattering phase throughout their orbits, we computed the contrast ratio from Equation 3 at each observable hour and used the resulting median contrast ratio ($\widetilde{F_{p}/F_{\star}}$) for our calculations. Tables 1 and 2 show the expected achievable post-processing contrasts from the instrument models in Section 2.2 for the GMT and ELT, respectively, at S/N = 5 for a one hour integration of the O2 A-band at $R=100,000$ for different host magnitudes and angular separations.
@@
To translate integration time to years of observations:
@@
where $Q=2$ for second quartile and $Q=4$ for first quartile observing conditions.
@@
Mag|$\theta$ 15 mas 30 mas 50 mas 100 mas 120 mas
1 5.85E-09 4.53E-09 3.51E-09 2.55E-09 2.36E-09
3 1.93E-08 1.61E-08 1.32E-08 1.02E-08 9.53E-09
5 8.01E-08 6.89E-08 5.85E-08 4.68E-08 4.41E-08
7 4.55E-07 4.02E-07 3.49E-07 2.88E-07 2.73E-07
9 4.31E-06 3.89E-06 3.45E-06 2.89E-06 2.73E-06
10 1.82E-05 1.66E-05 1.50E-05 1.26E-05 1.16E-05
@@
1 5.00E-09 4.07E-09 3.89E-09 3.11E-09 2.58E-09 1.97E-09
3 1.30E-08 1.06E-08 1.02E-08 8.33E-09 7.06E-09 5.50E-09
5 3.91E-08 3.21E-08 3.12E-08 2.64E-08 2.26E-08 1.75E-08
7 1.49E-07 1.22E-07 1.19E-07 1.02E-07 8.82E-08 6.85E-08
9 7.95E-07 6.54E-07 6.41E-07 5.60E-07 4.91E-07 3.90E-07
10 2.34E-06 1.93E-06 1.89E-06 1.67E-06 1.49E-06 1.20E-06
Table 2: ELT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system, 128 deformable mirror actuators, and Q2 observing conditions. For different AO control systems, actuators, and observing conditions see Table 4 in Appendix A.
@@ 3 Survey Simulation
In order to realistically simulate the execution of an observing survey of nearby EECs, we followed the survey simulation methodology of Hardegree-Ullman et al. (2023). First, we generated 1000 simulated local neighborhood ($d<20$ pc) exoplanet populations in Bioverse based on exoplanet demographics of Bergsten et al. (2022) and computed the time it will take to probe Earth-like O2 levels for each simulated EEC. We then sorted each population of simulated exoplanets by the number of years it would take to probe O2 and calculated the median value of the first through nth planets. The results are shown in Figure 1 for optimized instrument and observing modes, which are Max Speed, Q1 for the GMT, and Max Speed, Q2, 128 actuators for the ELT. An assessment of different instrument and observing modes is given in Appendix A.
Figure 1 shows that the ELT would be able to survey $\sim$3$\times$ more planets than the GMT for a survey of similar length, approximately commensurate with the $\sim$3$\times$ more collecting area. In a 10-year survey, if all EECs have Earth-like levels of O2, the GMT could probe up to $\sim$5 planets, and the ELT could probe up to $\sim$15. Note, that this simulation did not consider that these planets would likely have overlapping observing windows, so the number of planets surveyed for O2 is an upper limit. It is also important to highlight that over 85% of the simulated planets which could be probed within a 10-year survey across all survey simulations orbit M dwarfs, and about 75% of these simulated planets are within 13 pc of the Sun (Figure 2).
[…]
@@ 3.1 Super-Earths
Next, we carried out a simulation to explore the observability of O2 in all super-Earth exoplanets. The previous survey simulation focused on planets nearly the same size as Earth ($0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$). While not all exoplanets smaller than $1.4~{}R_{\oplus}$ will necessarily be Earth-like, a larger population including all super-Earth exoplanets would allow us to observationally constrain the limits of Earth-like planet atmospheres. In the following simulation, we assessed the ability to probe Earth-like O2 levels on Earth to super-Earth sized planets ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$), where the upper limit is guided by the exoplanet radius valley, Fulton et al., 2017; Hardegree-Ullman et al., 2020). Results from the super-Earth survey simulation are given in Figure 3. We found that about 1.5$\times$ more planets could be surveyed if the survey is broadened from Earth-sized to all Earth and super-Earth-sized planets.
[…]
@@ 3.2 Known Habitable Zone Exoplanets
[…]
@@ 4 Hypothesis Testing
In our test case, we aim to assess the statistical power to test the above hypothesis in a 10-year survey with the ELT considering different values of $f^{\mathrm{O}_{\mathrm{2}}}_{\mathrm{EEC}}$ (ten levels between 0.1 and 1) and $\eta_{\oplus}$ (five levels between 0.075 and 0.3). For simplicity, we set $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}=1\%$. We imposed log-uniform priors $a_{\mathrm{inner}}=[0.1,2]$, $\Delta a=[0.01,3]$, $f_{\mathrm{HZ}}=[0.01,1]$, $f_{\text{non-HZ}}/f_{\mathrm{HZ}}=[0.01,1.0]$.
@@ 5 Discussion
The detection of O2 in the atmospheres of Earth-analogs, if found, will no doubt have profound impacts on astronomy and astrobiology. Our simulations show that Earth-like O2 levels could be probed on EECs with the GMT and ELT with planned instrumentation. There are already a half-dozen known probable super-Earths which may be ideal candidates to start searching for oxygen (Section 3.2), and planets like Proxima Centauri b could be high-priority science targets for the GMT and ELT.
@@ 5.2 Other Atmospheric Biosignatures
Figure 7: The ability to probe H2O (blue dashed), O2 (green dashed dotted), CH4 (orange dotted), CO2 (pink dashed), and CO (black dotted) in the atmospheres of exoplanets in the habitable zones of G2 V through M8 V stars at $2\lambda/D$ for distances of 5 pc (left) and 10 pc (right). Lines are shown for a $\sim$6-meter JWST or anticipated HWO diameter aperture (orange solid), the GMT (blue solid), the TMT (green solid), and ELT (yellow solid). If a habitable zone region falls above the line for a telescope, the biosignature could be accessible to that facility at an inner working angle of $2\lambda/D$.
@@ 5.4 Thirty Meter Telescope
[…]
@@ 6 Summary and Conclusions
The Bioverse survey simulation and hypothesis testing framework enables comprehensive and quantitative assessments of the capabilities of upcoming ground-based telescopes and space missions to study exoplanets. We expanded the Bioverse framework to assess the ability of the ELT and the GMT to probe Earth-like levels of O2 on nearby ($<$20 pc) habitable zone Earth-sized exoplanets via extreme adaptive optics direct imaging and high resolution spectroscopy. The following list summarizes the main conclusions of this work:
[…]
@@
  • A 10-year survey with the ELT could determine if habitable zone exo-Earths are more likely to have O2. The diagnostic power of the survey to test this habitable zone oxygen hypothesis is sensitive to the fraction of EECs with Earth-like levels of O2 and the value of $\eta_{\oplus}$, but not very sensitive to maximum planet size (Figures 5 and 6).
@@
The upcoming 25–40-meter class telescopes are going to be the first facilities capable of robustly detecting biosignatures with extreme adaptive optics direct imaging and high resolution spectroscopy. This method is much more promising than transmission spectroscopy in the search for O2 (Hardegree-Ullman et al., 2023), and a handful of high-priority target planets are already available. The ability to validate biosignature detections and test the habitable zone oxygen hypothesis and other biosignature hypotheses, however, depends on probing a larger sample of planets than has been currently discovered.
@@ 7 Acknowledgments
We acknowledge the work of Bryan Wang to improve the Bioverse code. KKH-U acknowledges Megan Weiner Mansfield for general discussions about M dwarf exoplanet atmospheres and Suri Rukdee for discussions about the LTT 1445 A system.
@@ Appendix A Different Instrument Modes
For both the GMT and ELT, we first assessed different instrument setups and observing conditions. The GMagAO-X on GMT setup assumes a baseline number of 182 deformable mirror actuators, and we looked at two different AO control system optimization strategies: Max Speed, and Strehl Optimization. For PCS on the ELT, additional calculations were made for both a 128 or 200 actuator setup. Table 3 shows results for the different AO control systems and observing conditions for the GMT, and Table 4 shows results for different AO control systems, actuator counts, and observing conditions for the ELT.
@@
[…]
A higher resolution spectrograph produces narrower and deeper absorption and emission lines. For O2, higher resolution gives access to more relative system velocity space. We modeled an $R=500,000$ spectrograph coupled to our AO systems, for which severe line blending in the O2 A-band occurs when $|\Delta\mathrm{RV}|<10$ km s-1 and $40\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<48\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023). As noted by Fowler et al. (2023), post-processing efficiency does not change at $R>100,000$, so we are able to use Tables 3 and 4 for our $R=500,000$ calculations. The lower panel of Figure 8 shows the results for $R=500,000$ compared to $R=100,000$, with a $\sim$1 or 2 planet increase in planets surveyed for a survey of 10 years, suggesting a higher resolution spectrograph is not necessarily a good investment for this science case.

### v2

@@ 2.3 Time Needed to Probe O2
With our new instrument model, we now have to determine how long it will take to probe O2. For each simulated exoplanet, we first computed the total number of hours the planet would be observable from the GMT and ELT locations for one year of observations ($N_{\text{hr-obs}}$) accounting for the aforementioned ground-based observing limitations and exoplanet orbital properties. Since the modeled planets change illumination and scattering phase throughout their orbits, we computed the contrast ratio from Equation 3 at each observable hour and used the resulting median contrast ratio ($\widetilde{F_{p}/F_{\star}}$) for our calculations. Tables 3 and 4 show the expected achievable post-processing contrasts from the instrument models in Section 2.2 for the GMT and ELT, respectively, at S/N = 5 for a one hour integration of the O2 A-band at $R=100,000$ for different host magnitudes and angular separations.
@@
To translate total integration time to years of observations, we need to account for both target observability ($N_{\text{hr-obs}}$) and acceptable site observing conditions ($Q$):
@@
where $Q=4$ for first quartile observing conditions (best observing site weather, ideal for extreme-AO, $\sim$25of telescope observing time, will take four times longer to achieve S/N) and $Q=2$ for second quartile observing conditions (best to average observing site weather, $\sim$50of telescope observing time, will take twice as long to achieve S/N). It is worth emphasizing that $N_{\text{hr-obs}}$ only counts times at which the relative system velocities yield unblended lines and the planet is observable from the observing site. We simulate first and second quartile observing conditions for each instrument mode, and assess the trade-offs of different observing scenarios in the Appendix.
@@
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 0.57 3.43 3.63 3.78 3.84 3.84
3 0.57 3.40 3.60 3.74 3.80 3.81
5 0.55 3.29 3.48 3.63 3.68 3.69
7 0.48 2.90 3.07 3.19 3.24 3.25
9 0.30 1.78 1.88 1.96 1.99 1.99
10 0.16 0.94 1.00 1.04 1.06 1.06
Table 2: End-to-end GMagAO-X planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system and Q1 observing conditions. For different AO control systems and observing conditions see Table 5 in Appendix A.
GMT, Max Speed, Q1
Mag|$\theta$ 5 mas 10 mas 15 mas 30 mas 50 mas 120 mas
1 3.06E-08 6.09E-09 5.85E-09 4.53E-09 3.51E-09 2.35E-09
3 9.79E-08 1.97E-08 1.93E-08 1.61E-08 1.32E-08 9.50E-09
5 4.01E-07 8.10E-08 8.00E-08 6.88E-08 5.83E-08 4.39E-08
7 2.26E-06 4.57E-07 4.54E-07 4.00E-07 3.47E-07 2.71E-07
9 2.12E-05 4.30E-06 4.30E-06 3.87E-06 3.43E-06 2.72E-06
10 8.87E-05 1.80E-05 1.81E-05 1.65E-05 1.49E-05 1.15E-05
@@
1 2.55E-09 2.08E-09 1.99E-09 1.55E-09 1.25E-09 9.36E-10
3 7.12E-09 5.84E-09 5.64E-09 4.64E-09 3.90E-09 2.95E-09
5 2.48E-08 2.04E-08 1.99E-08 1.68E-08 1.43E-08 1.09E-08
7 1.03E-07 8.49E-08 8.30E-08 7.14E-08 6.17E-08 4.78E-08
9 5.72E-07 4.72E-07 4.64E-07 4.06E-07 3.57E-07 2.82E-07
10 1.62E-06 1.34E-06 1.32E-06 1.17E-06 1.03E-06 8.24E-07
Table 4: ELT achievable $F_{p}/F_{\star}$ contrast at S/N=5 for 1 hour integration times using high contrast imaging with high-resolution spectroscopy for different host star magnitudes and planet-star angular separations ($\theta$). This table is for a Max Speed AO control system, 128 deformable mirror actuators, and Q1 observing conditions. For different AO control systems, actuators, and observing conditions see Table 7 in Appendix A.
@@ 3 Survey Simulation
In order to realistically simulate the execution of an observing survey of nearby EECs, we followed the survey simulation methodology of Hardegree-Ullman et al. (2023). First, we generated 1000 simulated local neighborhood ($d<20$ pc) exoplanet populations in Bioverse based on exoplanet demographics of Bergsten et al. (2022) and computed the time it will take to probe Earth-like O2 levels for each simulated EEC. We then sorted each population of simulated exoplanets by the number of years it would take to probe O2 and calculated the median value of the first through nth planets. The results are shown in Figure 4 for optimized instrument and observing modes, which are Max Speed, Q1 for the GMT, and Max Speed, Q1, 128 actuators for the ELT. An assessment of different instrument and observing modes is given in Appendix A.
Figure 4 shows that the ELT would be able to survey $\sim$3$\times$ more planets than the GMT for a survey of similar length, approximately commensurate with the $\sim$3$\times$ more collecting area. For example, in a 10-year survey from start to finish, if all EECs have Earth-like levels of O2, the GMT could probe up to $\sim$7 planets, and the ELT could probe up to $\sim$19. Note, that this simulation did not consider that these planets would likely have overlapping observing windows, so the number of planets surveyed for O2 is an upper limit. Further observing optimization (e.g., observing exoplanets near maximum phase, or combining measurements from multiple telescopes) should be possible, but was not explored in this study. It is also important to highlight that over 85% of the simulated planets which could be probed within a 10-year survey across all survey simulations orbit M dwarfs, and about 75% of these simulated planets are within 13 pc of the Sun (Figure 5).
Figure 4: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 on exo-Earth candidates within 20 pc with the ELT (blue dots) and the GMT (orange stars). A survey of the same duration on both telescopes would allow the ELT to probe about three times more planets for O2 than the GMT. The large error bars are driven by the wide range of possible orbital parameters for our simulated exoplanets.
@@ 3.1 Super-Earths
Next, we carried out a simulation to explore the observability of O2 in all super-Earth exoplanets. The previous survey simulation focused on planets nearly the same size as Earth ($0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$). While not all exoplanets smaller than $1.4~{}R_{\oplus}$ will necessarily be Earth-like, a larger population including all super-Earth exoplanets would allow us to observationally constrain the limits of Earth-like planet atmospheres. In the following simulation, we assessed the ability to probe Earth-like O2 levels on Earth to super-Earth sized planets ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$), where the upper limit is guided by the exoplanet radius valley (Fulton et al., 2017; Hardegree-Ullman et al., 2020). Results from the super-Earth survey simulation are given in Figure 6. We found that about 1.5$\times$ more planets could be surveyed if the survey is broadened from Earth-sized to all Earth and super-Earth-sized planets.
Figure 6: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 on habitable zone Earth and super-Earth-sized candidates ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$) within 20 pc with the ELT (blue dots) and the GMT (orange stars). Similar to the exo-Earth simulation in Figure 4, the ELT could probe about three times more planets than the GMT. Overall, a super-Earth survey could probe about 1.5 times more planets than a survey of planets up to 1.4 $R_{\oplus}$.
@@ 3.2 Known Habitable Zone Exoplanets
Figure 7: Simulation results for a survey of Earth-like concentrations of O2 at S/N=5 for five known, and one candidate (LTT 1445 A d) habitable zone exoplanets within 20 pc with the ELT (blue) and the GMT (orange). The ELT and GMT should both be able to probe for O2 on Proxima Centauri b in less than a month. LTT 1445 A d is the only planet in the simulation for which Earth-like levels of O2 could not be probed with the GMT, but it could be probed with the ELT. More precise planet and orbital parameters would minimize the uncertainties on these observing simulations. The minimum mass ($M_{\oplus}\sin i$) of each planet is given above each relatively-sized point (Faria et al., 2022; Bonfils et al., 2018; Astudillo-Defru et al., 2017; Lavie et al., 2023; Anglada-Escudé et al., 2013).
@@ 4 Hypothesis Testing
In our test case, we aim to assess the statistical power to test the above hypothesis in a 10-year survey with the ELT considering different values of $f^{\mathrm{O}_{\mathrm{2}}}_{\mathrm{EEC}}$ (ten levels between 0.1 and 1) and $\eta_{\oplus}$ (five levels between 0.075 and 0.3). For simplicity, we set $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}=1\%$. A higher value of $f^{\mathrm{O}_{\mathrm{2}}}_{\text{non-EEC}}$ would inject more false-positive O2 signals and skew our results to require fewer detections to test our hypothesis. We intentionally keep this value low because future surveys, like the one used in this hypothesis test, will likely only probe planets within the habitable zone defined either by incident stellar flux levels, or confirmed with liquid water detections on the habitable zone planets. We imposed log-uniform priors $a_{\mathrm{inner}}=[0.1,2]$, $\Delta a=[0.01,3]$, $f_{\mathrm{HZ}}=[0.01,1]$, $f_{\text{non-HZ}}/f_{\mathrm{HZ}}=[0.01,1.0]$.
@@ 5 Discussion
The detection of O2 in the atmospheres of Earth-analogs, if found, will no doubt have profound impacts on astronomy and astrobiology. Our simulations show that Earth-like O2 levels could be probed on EECs with the GMT and ELT with planned instrumentation. There are already a half-dozen known probable super-Earths which may be ideal candidates to start searching for oxygen (Section 3.2), and planets like Proxima Centauri b could be high-priority science targets for the GMT and ELT. Below we discuss simulation limitations and survey feasibility, O2  other biosignatures, and the importance of population-level biosignature interpretation, and end with a brief discussion of the TMT.
5.1 Limitations and Survey Feasibility
Our simulations are based on loosely constrained (e.g., $\eta_{\oplus}$) or unknown astrophysical parameters (e.g., orbital parameters). These contribute significantly to the uncertainties of our survey simulations. Exoplanet occurrence rate studies with TESS, Roman, and PLATO, and radial velocity surveys to find new EECs around nearby stars could help further reduce survey uncertainties and find new candidates for biosignature searches.
@@ 5.1 Limitations and Survey Feasibility
Since the extreme-AO systems for the ELTs are still in their planning stages, we rely on instrument parameters that are not yet finalized or that are based on current systems. Consequentially, our simulations can also be used to provide guidance to instrument teams as to the requirements for biosignature searches with the ELTs.
In Section 3, we did not consider that multiple target planets could be observable at the same time, or by multiple ELTs. We also used the median calculated contrast ratio for each simulated planet (using only contrast ratios for times the planets were observable in one year of simulated observations). For simulation purposes, these are reasonable assumptions to provide upper limits to yields for an exoplanet biosignature survey for individual telescopes.
Our 10-year survey assumed we would observe during Q1 observing conditions. Since the ELTs are not solely dedicated to exoplanet science, it is unlikely that all Q1 observing time would be dedicated to the search for biosignatures. Though, exoplanets are not always observable or at ideal phases for direct imaging during Q1 observing conditions. As we assess in Appendix A, it makes little difference whether a survey is conducted in Q1 or Q2 observing conditions, which would likely be a factor in designing a biosignature survey. Until additional nearby EECs are identified, we can only report upper limits to survey yields for a hypothetical observing scenario as presented here.
@@ 5.3 Other Atmospheric Biosignatures
Figure 10: The inner working angles at $2\lambda/D$ for 0.6–2 $\mu$m are plotted for a $\sim$6-meter JWST or anticipated HWO diameter aperture (orange solid line), the GMT (blue solid line), the TMT (green solid line), and the ELT (yellow solid line). The wavelengths of H2O (blue dashed), O2 (green dashed dotted), CH4 (orange dotted), CO2 (pink dashed), and CO (black dotted) for this wavelength region are plotted as vertical lines. At distances of 5 pc (left) and 10 pc (right), The shaded horizontal regions show the conservative habitable zones for representative G2 V through M8 V stars. If a habitable zone region falls above the inner working angle of a telescope, an exoplanet and its biosignatures should be observable at that wavelength range. For example, below 1 $\mu$m, the GMT should be able to observe O2, CH4, and H2O for planets in the habitable zones of a hypothetical M4 V star at 5 pc. The GMT will only be able to access biosignatures between 1 and 2 $\mu$m if planets fall in certain regions of the habitable zone of this hypothetical star, and it will not be able to access biosignatures beyond 2 $\mu$m.
@@ 5.5 Thirty Meter Telescope
Having multiple facilities capable of searching for biosignatures is advantageous. When we are looking for S/N=5 biosignature detections over years of observations and hundreds of datasets, it will be important to have independent verification from, e.g., different telescopes and processing techniques to minimize false-positive scenarios. Multiple facilities can both expedite data collection on individual targets and potentially expand the target pool amenable to biosignature surveys if data from different instruments can be combined, and since weather conditions at one facility become less of an observing bottleneck. Hoeijmakers et al. (2018) successfully demonstrated combining more than 2,000 high-resolution reflected light spectra from different 2–8-meter class telescopes to measure the albedo of the hot Jupiter $\tau$ Boötis b and achieved a planet-to-star contrast of $\sim$10-5. The GMT, TMT, and ELT will be able to build upon these methods to reach smaller planets and potentially detect biosignatures.
@@ 6 Summary and Conclusions
The Bioverse survey simulation and hypothesis testing framework enables comprehensive and quantitative assessments of the capabilities of upcoming ground-based telescopes and space missions to study exoplanets. We expanded the Bioverse framework to assess the ability of the ELT and the GMT to probe Earth-like levels of O2 on nearby ($<$20 pc) habitable zone Earth-sized exoplanets via extreme-AO direct imaging and high resolution spectroscopy. The following list summarizes the main conclusions of this work:
  • Accounting for recent planet occurrence rate calculations, relative system velocities, and realistic target observability, we conducted a survey simulation to determine how many exo-Earths ($0.8S^{0.25}<R_{p}<1.4~{}R_{\oplus}$) and super-Earths ($0.8S^{0.25}<R_{p}<1.8~{}R_{\oplus}$) could be probed for Earth-like levels of O2 at S/N=5 with the GMT and ELT. In 10 years, the GMT could survey up to $\sim$7 and $\sim$10 Earths and super-Earths, respectively, while the ELT could survey up to $\sim$19 and $\sim$31 (Figures 4 and 6).
@@
  • A 10-year survey with the ELT could determine if habitable zone exo-Earths are more likely to have O2. The diagnostic power of the survey to test this habitable zone oxygen hypothesis is sensitive to the fraction of EECs with Earth-like levels of O2 and the value of $\eta_{\oplus}$, but not very sensitive to maximum planet size (Figures 8 and 9).
@@
The upcoming 25–40-meter class telescopes are going to be the first facilities capable of robustly detecting biosignatures with extreme-AO direct imaging and high resolution spectroscopy. This method is much more promising than transmission spectroscopy in the search for O2 (Hardegree-Ullman et al., 2023), and a handful of high-priority target planets are already available. The ability to validate biosignature detections and test the habitable zone oxygen hypothesis and other biosignature hypotheses, however, depends on probing a larger sample of planets than has been currently discovered.
@@ 7 Acknowledgments
We acknowledge the work of Bryan Wang to improve the Bioverse code. KKH-U acknowledges Megan Weiner Mansfield for general discussions about M dwarf exoplanet atmospheres and Suri Rukdee for discussions about the LTT 1445 A system. We thank the anonymous referee for their helpful comments which significantly improved the clarity of this manuscript.
@@ Appendix A Different Instrument Modes
For both the GMT and ELT, we first assessed different instrument setups and observing conditions. The GMagAO-X on GMT setup assumes a baseline number of 188 deformable mirror actuators, and we looked at two different AO control system optimization strategies: Max Speed, and Strehl Optimization. For PCS on the ELT, additional calculations were made for both a 128 or 200 actuator setup. Table 5 shows results for the different AO control systems and observing conditions for the GMT, and Table 7 shows results for different AO control systems, actuator counts, and observing conditions for the ELT. For each of these setups, we also give the end-to-end planet throughput for different host star magnitudes and angular separations for the GMT/GMagAO-X in Table 6 and the ELT/PCS in Table 8.
@@
Max Speed, Q1 Max Speed, Q2
1 0.57 3.43 3.63 3.78 3.84 3.84 0.56 3.35 3.55 3.69 3.75 3.75
3 0.57 3.40 3.60 3.74 3.80 3.81 0.55 3.30 3.50 3.64 3.69 3.70
5 0.55 3.29 3.48 3.63 3.68 3.69 0.52 3.14 3.32 3.46 3.51 3.52
7 0.48 2.90 3.07 3.19 3.24 3.25 0.43 2.59 2.74 2.85 2.89 2.90
9 0.30 1.78 1.88 1.96 1.99 1.99 0.21 1.24 1.31 1.36 1.38 1.39
10 0.16 0.94 1.00 1.04 1.06 1.06 0.08 0.48 0.51 0.53 0.53 0.54
Mag Strehl Optimization, Q1 Strehl Optimization, Q2
1 0.57 3.43 3.63 3.78 3.84 3.84 0.56 3.35 3.55 3.69 3.75 3.75
3 0.57 3.41 3.61 3.76 3.82 3.82 0.55 3.31 3.50 3.65 3.70 3.71
5 0.56 3.36 3.55 3.70 3.76 3.76 0.54 3.22 3.41 3.55 3.60 3.61
7 0.54 3.23 3.42 3.56 3.62 3.62 0.50 2.99 3.17 3.29 3.35 3.35
9 0.49 2.92 3.09 3.22 3.27 3.27 0.41 2.47 2.62 2.73 2.77 2.77
10 0.44 2.64 2.80 2.91 2.96 2.96 0.34 2.06 2.18 2.26 2.30 2.30
Table 6: End-to-end GMagAO-X planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization) and Q1 and Q2 observing conditions.
@@
Max Speed, 128 actuators, Q1 Max Speed, 128 actuators, Q2
1 2.74 3.88 3.92 3.95 3.96 3.96 2.32 3.29 3.33 3.36 3.36 3.36
3 2.73 3.87 3.92 3.95 3.95 3.95 2.32 3.28 3.32 3.35 3.35 3.35
5 2.72 3.86 3.90 3.94 3.94 3.94 2.30 3.26 3.30 3.33 3.33 3.34
7 2.69 3.80 3.85 3.88 3.88 3.89 2.26 3.20 3.24 3.26 3.27 3.27
9 2.48 3.51 3.55 3.59 3.59 3.59 2.00 2.84 2.87 2.90 2.90 2.90
10 2.13 3.01 3.05 3.07 3.08 3.08 1.57 2.22 2.25 2.27 2.27 2.27
Mag Max Speed, 200 actuators, Q1 Max Speed, 200 actuators, Q2
1 2.94 4.16 4.21 4.25 4.25 4.25 2.74 3.88 3.93 3.96 3.97 3.97
3 2.93 4.15 4.20 4.24 4.24 4.24 2.73 3.87 3.92 3.95 3.95 3.95
5 2.91 4.12 4.17 4.21 4.21 4.21 2.70 3.83 3.88 3.91 3.91 3.91
7 2.82 4.00 4.04 4.08 4.08 4.08 2.58 3.65 3.70 3.73 3.73 3.73
9 2.37 3.36 3.40 3.43 3.43 3.43 1.92 2.73 2.76 2.78 2.79 2.79
10 1.88 2.66 2.69 2.71 2.72 2.72 1.26 1.78 1.80 1.82 1.82 1.82
1 2.74 3.88 3.92 3.95 3.96 3.96 2.32 3.29 3.33 3.36 3.36 3.36
3 2.73 3.87 3.92 3.95 3.95 3.95 2.32 3.28 3.32 3.35 3.35 3.35
5 2.73 3.86 3.91 3.94 3.94 3.95 2.31 3.27 3.31 3.34 3.34 3.34
7 2.71 3.84 3.89 3.92 3.92 3.92 2.28 3.24 3.27 3.30 3.30 3.31
9 2.67 3.78 3.82 3.86 3.86 3.86 2.22 3.14 3.18 3.20 3.21 3.21
10 2.62 3.71 3.76 3.79 3.79 3.79 2.14 3.04 3.07 3.10 3.10 3.10
1 2.94 4.16 4.21 4.25 4.25 4.25 2.74 3.88 3.93 3.96 3.97 3.97
3 2.93 4.15 4.20 4.24 4.24 4.24 2.73 3.87 3.92 3.95 3.96 3.96
5 2.92 4.14 4.19 4.22 4.23 4.23 2.71 3.85 3.89 3.92 3.93 3.93
7 2.89 4.09 4.14 4.18 4.18 4.18 2.66 3.77 3.81 3.85 3.85 3.85
9 2.80 3.97 4.02 4.06 4.06 4.06 2.51 3.56 3.60 3.64 3.64 3.64
10 2.71 3.84 3.88 3.92 3.92 3.92 2.35 3.33 3.37 3.40 3.40 3.40
Table 8: End-to-end ELT/PCS planet throughput in percentages (from 0% to 100%) for different host star magnitudes and planet-star angular separations ($\theta$). Tables are given for different AO control systems (Max Speed and Strehl Optimization), 128 and 200 deformable mirror actuators, and Q1 and Q2 observing conditions.
Figure 11: Simulation results for a survey of Earth-like O2 levels at S/N=5 on exo-Earth candidates within 20 pc with different AO control systems and observing conditions for the GMT (upper left), additionally accounting for different deformable mirror actuator counts on the ELT (upper right), and considering an optimized setup for both the GMT and ELT but assessing the differences between R=100,000 and R=500,000 (lower panel).
A higher resolution spectrograph produces narrower and deeper absorption and emission lines. For O2, higher resolution gives access to more relative system velocity space. We modeled an $R=500,000$ spectrograph coupled to our AO systems, for which severe line blending in the O2 A-band occurs when $|\Delta\mathrm{RV}|<10$ km s-1 and $40\ \mathrm{km\ s}^{-1}<|\Delta\mathrm{RV}|<48\ \mathrm{km\ s}^{-1}$ (López-Morales et al., 2019; Hardegree-Ullman et al., 2023). As noted by Fowler et al. (2023), post-processing efficiency does not change at $R>100,000$, so we are able to use Tables 5 and 7 for our $R=500,000$ calculations. The lower panel of Figure 11 shows the results for $R=500,000$ compared to $R=100,000$, with a $\sim$1 or 2 planet increase in planets surveyed for a survey of 10 years, suggesting a higher resolution spectrograph is not necessarily a good investment for this science case.